OpenStudy (anonymous):

please help! sin^2(x)+sinx=0 find all solutions in the interval [0,2pi)

OpenStudy (anonymous):

I know you are supposed to break it down into \[\sin(\sin+1)=0\] so then it goes to \[\sin=0 \] and \[\sin+1=0\] but I don't know what to do next

OpenStudy (nikato):

 solve for x now